OpenStudy (anonymous):

I need to find the value of k such that the integral from -infinity to infinity of f(x)dx = 1. How can I do this with the function f(x) = k |x|/(x^2 +4) ?

OpenStudy (anonymous):

I have tried to integrate to separate out k, but it's not working really well.

OpenStudy (anonymous):

hello!

OpenStudy (helder_edwin):

find k such that \[ \large 1=\int_{-\infty}^{+\infty}\frac{k|x|}{x^2+4}\,dx \] ??

OpenStudy (anonymous):

That's what I tried to do, but I don't know how work out the last bits, where I have an exponent on either side and I'm trying to get k.

OpenStudy (helder_edwin):

the function inside the integral is symetric with respect to y-axis. did u notice that?

OpenStudy (anonymous):

So I get\[1= 1/2 k \int\limits_{?}^{?}1/(x^2 +4)\]

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

oh, I see.

OpenStudy (anonymous):

well, I don't know actually, since the variable is x?

OpenStudy (helder_edwin):

\[ \large \frac{|-x|}{(-x)^2+4}=\frac{|x|}{x^2+4} \] right?

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

so \[ \large \int_{-\infty}^{+\infty}\frac{k|x|}{x^2+4}\,dx= 2\int_{0}^{+\infty}\frac{kx}{x^2+4}\,dx \] \[ \large =2\lim_{b\to+\infty}k\int_0^b\frac{x}{x^2+4}\,dx =2k\lim_{b\to+\infty}\int_0^b\frac{x}{x^2+4}\,dx \]

OpenStudy (helder_edwin):

putting u=x^2+4 then du=2x dx then x dx=du/2

OpenStudy (anonymous):

I've gotten\[1= (1/2) k \int\limits_{-\infty}^{\infty} 1/(x^2 +4) dx\] down to \[1= (1/2) k ((1/2) \tan^-1(x/2) or something like that\]

OpenStudy (anonymous):

but wait, you still have a x at the top when you took out the 2 from the intergral

OpenStudy (helder_edwin):

\[ \large =2k\lim_{b\to+\infty}\int_0^b\frac{1}{u}\,\frac{du}{2}= \frac{2k}{2}\lim_{b\to+\infty}\int_0^b\frac{du}{u} \]

OpenStudy (anonymous):

What you last typed is what I have, sort of. I don't understand the 0 to b thing, but I do understand the du part.

OpenStudy (anonymous):

and 0 to b covers -infinity to infinity?

OpenStudy (helder_edwin):

yes. since the function u r integrating is symetric with respect to y-axis, the area between the curve and the x-axis is the same to both sides of the y-axis

OpenStudy (anonymous):

I have a question: how did you get the 2 outside the integral right before the rest of your work? When I did it, I went right to du and ended up with a (1/2) outside the integral. How did you avoid that?

OpenStudy (helder_edwin):

the total area is twice the area either to the left or the right.

OpenStudy (anonymous):

wait, since going from 0 to infinity can be the same as going from 0 to -infinity (same infinite length, I can 'split' the integral? As if I will multiply by two to get the final answer?

OpenStudy (helder_edwin):

yes. it should work if the integral converges.

OpenStudy (anonymous):

So, now I am at a similar spot. I know the arc tan trig identity work here for du/u, but how can I solve this out for k?

OpenStudy (helder_edwin):

let mi finish. ok?

OpenStudy (anonymous):

okay :D

OpenStudy (helder_edwin):

\[ \large =k\lim_{b\to+\infty}\left(\int_0^1\frac{du}{u}+\int_1^b\frac{du}{u}\right) \] \[ \large =k\left[\lim_{a\to 0+}\int_a^1\frac{du}{u}+ \lim_{b\to+\infty}\int_1^b\frac{du}{u}\right] \]

OpenStudy (helder_edwin):

\[ \large =k\left[\lim_{a\to0+}(\ln1-\ln a)+ \lim_{b\to+\infty}(\ln b-\ln 1)\right] =k[-(-\infty)+(+\infty)] \]

OpenStudy (helder_edwin):

done.

OpenStudy (anonymous):

wait...k equals...infinity? I'm not sure I understand...,

OpenStudy (helder_edwin):

me neither

OpenStudy (helder_edwin):

r u sure about the |x| in the numerator.

OpenStudy (anonymous):

ur, no, if it's 1= k[inifinity + infinity], does k = 1/infinity= facepalm?

OpenStudy (anonymous):

yeah, function reads f(x) = k|x|/((x^2) +4)

OpenStudy (anonymous):

I'm supposed to solve it so that if it was evaluated from -infinity to infinity, it would equal 1, depending on the correct value of k

OpenStudy (helder_edwin):

\[ \large =k[+\infty+(+\infty)]=k(+\infty) \]

OpenStudy (helder_edwin):

\[ \large =+\infty \]

OpenStudy (helder_edwin):

it is impossible. all the work shows the integral diverges.

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

what if the |x| wasn't in the problem? Would that change things?

OpenStudy (anonymous):

b/c I think they don't give impossible problems on the first day of school, lol

OpenStudy (helder_edwin):

then we would have \[ \large 1=2k\int_0^{+\infty}\frac{1}{x^2+4}\,dx= 2k\lim_{b\to+\infty}\int_0^b\frac{dx}{x^2+4} \] \[ \large =2k\lim_{b\to+\infty}\frac{1}{2}\arctan\frac{x}{2}\biggr|_0^b \]

OpenStudy (anonymous):

yeah, that's what I have

OpenStudy (anonymous):

but any way you turn it, you are going to end up multiply k by zero, unless i'm going it wrong

OpenStudy (helder_edwin):

\[ \large 2k\frac{1}{2}\lim_{b\to+\infty}\arctan\frac{b}{2}=k\frac{\pi}{2} \]

OpenStudy (helder_edwin):

that is \[ \large 1=k\frac{\pi}{2} \] then \[ \large k=\frac{2}{\pi} \]

OpenStudy (anonymous):

How does arctan(b/2) (input infinity) become...oh

OpenStudy (anonymous):

no, what?

OpenStudy (helder_edwin):

\[ \large \lim_{x\to+\infty}\arctan x=\frac{\pi}{2} \]

OpenStudy (anonymous):

that is true

OpenStudy (anonymous):

where x = b/2?

OpenStudy (helder_edwin):

well if \(b\to+\infty\) then \(b/2\to+\infty\) as well

OpenStudy (anonymous):

wait, you have that extra 2 by k again. I thought we got rid of that?

OpenStudy (helder_edwin):

we did!

OpenStudy (anonymous):

...so how did it reappear to reduce the other 1/2 from the trig identity? Maybe I missed something.

OpenStudy (anonymous):

We had k *lim(b-->+infinity) (du/u), and then we know that du/u = trig identity (1/2) arctan(x/2).

OpenStudy (helder_edwin):

u r forgetting this \[ \large \int_{-\infty}^{+\infty}\frac{1}{x^2+4}\,dx=2\int_0^{+\infty} \frac{1}{x^2+4}\,dx \]

OpenStudy (anonymous):

yeah, that's the part I'm talking about: we take du and that 2 cancels (2k/2)

OpenStudy (anonymous):

so we end up with k(integral) 1/(x^2 +4)

OpenStudy (anonymous):

how do we mysteriously get another 2?

OpenStudy (anonymous):

Sorry, I really want to understand this. I've got two more problems just like it.

OpenStudy (helder_edwin):